OpenStudy (anonymous):

graph the equation by plotting three points-3y=5x-7

OpenStudy (anonymous):

can someone please help me

OpenStudy (jdoe0001):

so plot 3 points :/

OpenStudy (anonymous):

that the thing I don't know what three points to use for x

OpenStudy (phi):

if you had the 3 points, could you plot them ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

the first thing I would do is divide both sides by -3 to get \[ y =\frac{5x-7}{-3} \]

OpenStudy (jdoe0001):

pick any value for "x" any value whatsoever

OpenStudy (anonymous):

like 1,0,-1 jdoe001

OpenStudy (jdoe0001):

sure

OpenStudy (anonymous):

that's wrong

OpenStudy (phi):

if you pick x= 1 that means replace x with 1 in the equation \[ y =\frac{5x-7}{-3} \\ y =\frac{5\cdot 1-7}{-3} \] now simplify: 5 times 1 is 5 5-7 is -2 you get \[ y = \frac{-2}{-3} = \frac{2}{3} \] that means your first point is (1, 2/3) not exactly pretty, because you have to guess 2/3 of a step for the y-value

OpenStudy (phi):

can you plot (1, 2/3) ?

OpenStudy (anonymous):

I can plot 1 but not 2/3 it has to be a whole number

OpenStudy (jdoe0001):

hehe, the "x" is so-called "independent variable" why is it called independent? because it can independently take whatever value it pleases

OpenStudy (phi):

try to "eyeball" it. divide the step into 3rds, and go up 2/3 res

OpenStudy (phi):